HomePage

News

Account

Network

Services

Login

en it

edit SideBar

Recent Changes
Printable View
Page History
Edit Page

Welcome to the CNR-IOM IT services homepage.

On the left sidebar you can see a selection of topics covering the account request, the computer connection and printer setup.

On the bottom left you can also choose the language in which you want to read the documentation.

Contact:

Stefano Bigaran +39 040 375 6424
Matteo Brutti +39 040 375 6460
Vanja Cvelbar +39 040 375 6456

Last News:

20170416-19 Network outage

EVENTS

  • Sunday 16th April at 6:02 CEST the link to and from Elettra and ISMAR Q2 was interrupted.
  • Monday 17th April at 22:48 CEST went down also the link to the GARR router that connects to the Internet the IOM headquarters and the part of ISMAR in Q2.
  • Tuesday 18th April the technicians available and the Elettra personnel managed to diagnose the problem finding that 5/8 optical fibres were severed. Some of them went dark during the diagnosis. This was most probably caused by a rodent chewing through the cable.
  • 10.32 CEST the Netgear switch went down on the rack with the following services: DNS, DHCP, VPN, Captive portal WiFi, Presenze site, Remote site, IT site, that were not accessible anymore. In the evening the connection to the GARR network was running again, but Elettra nad ISMAR were still unreachable. The machines that needed the above mentioned services to connect were obviously still disconnected.
  • Wednesday 19th April at 08:56 CEST the switch was replaced and reconfigured and the services were available again, allowing the full connection from the IOM HQ. Elettra, ISMAR Q2 and the WiFi networks were still unreachable.
  • 10:30 CEST meeting with the Elettra IT to plan the needed actions. The only available option is to let all the traffic flow through the only available fibre optics pair and to route it using the VLANs. The procedure to lay the new fibres is immediately started. The fibres are physically reconfigured and the routing is changed on the switches
  • 11:29 CEST the network is down to allow the reconfiguration. The operation is successfully ended at 13:10 CEST.
  • 13:30 CEST all the services including the WiFi network are available.

CRITICAL POINTS

  • At present the network is working, but it's very delicate because there is only one pair of fibres for both the internal and external link.
  • In the other server rack there is another Netgear switch of the same type.

ACTIONS

  • Replace the Netgear switch.
  • Harden the setup by moving one of the DHCP nodes to the other rack.
  • Install new fibre with anti-rodent protection.
  • Third party alert/news system
Page last modified on 2016-02-25 09:02 UTC
Istituto Officina dei Materiali - CNR P.IVA IT02118311006